Suppose that X1, X2, ..., Xn is an iid sample, each with probability p of being distributed as uniform over (-1/2,1/2) and with probability 1 - p of being distributed as uniform over (a) Find the cumulative distribution function (cdf) and the probability density function (pdf) of X1 (b) Find the maximum likelihood estimator (MLE) of p. c) Find another estimator of p using the method of moments (MOM)
Answer #1

The pdf of Xi is X, (n) = p × P) × = 1 if 0 < x1 < 1/2 1-0 0 otherwise The cdf of Xi is Fx,41) = P(X1 < x1) = 0 if x1 -1/2 0

(b) Write fx,(xi)-σ(u.)py, (1-p) where. σ(uī) = l if-1/2 < xi < 1 = 0 otherwise y. = 1.if xīE (-1/2. 0] = 0 otherwise = 0 ot

1/2 or, E (2-4X1)=p Now, E (2-4X)= p where, X = nz Hence MOM estimator of p=2-4X